OpenStudy (mikezack123):

If f(x) = x2^2 - 1, find f(3).

OpenStudy (mikezack123):

@satellite73 @Mertsj @phi

OpenStudy (anonymous):

it is 2x or x2

OpenStudy (anonymous):

or x^2^2

OpenStudy (anonymous):

is 2x

OpenStudy (nory):

f(x) = 2x^2 - 1 f(3) = 2*(3^2) - 1. What is 3 ^2? What is 2*3^2? What is 2*3^2 -1?

OpenStudy (mikezack123):

A. 5 B. 2 C. 8 D. -2

OpenStudy (mikezack123):

i got 17.... ?

OpenStudy (nory):

That's right... unless by x2^2 you meant (2x)^2 instead of 2x^2. But I think you got it right. Good job!

OpenStudy (mikezack123):

but these are the answers... A. 5 B. 2 C. 8 D. -2

OpenStudy (mikezack123):

:(

OpenStudy (nory):

That's weird. Draw a picture of the equation exactly how you remember it.

OpenStudy (mikezack123):

oops sorry heres the equation ,

OpenStudy (mikezack123):

If f(x) = x^2 - 1, find f(3).

OpenStudy (mikezack123):

either way you explained how to do it, thanks :) i got 8